# Dual Graph Attention based Disentanglement Multiple Instance Learning for Brain Age Estimation

### v1

@@ 4.1 Datasets and Preprocessing
A standard preprocessing pipeline using FMRIB Software Library (FSL, https://fsl.fmrib. ox.ac.uk/Woolrich et al. (2009); Smith et al. (2004); Jenkinson et al. (2012) is applied to all sMRI data. It includes several steps such as nonlinear registration Andersson et al. (2007) to the MNI152 standard space, brain extraction Smith (2002); Jenkinson et al. (2005), and normalization of voxel values. To ensure consistency of input sizes across different datasets, all pre-processed sMRI images are cropped based on a generated mask that removes uninformative zero-value pixels from the background. As a result, all MRIs have a voxel size of $160\times 192\times 160$ and an isotropic spatial resolution of 1 $mm^{3}$.
@@ 4.3 Network Training
To train the proposed DGA-DMIL network, we use ADAM optimizerKingma and Ba (2014) for optimization, and initialize it with a learning rate of 0.0001. The batch size is set to be 32 and the disentanglement weights agree $\lambda_{2}=0.1$, $\lambda_{3}=0.05$, and $\lambda_{4}=1$, based on their performance on the validation set. The weights in the CNN layers in the decoupling unit are initialized using the Kaiming weight initialization scheme He et al. (2015), and the weights in the fully connected layers are initialized with a normal distribution. The bias of the final fully connected layer is set equal to the average age of the training dataset for faster convergence. During training, the sum of the mean squared error loss and the decoupling loss is adopted as the total loss, and backpropagation is performed accordingly. All brain age estimation models are implemented in PyTorch.
@@
Fig. 6: Scatter diagrams of the estimated brain ages and chronological ages by different estimation models. (A) TSAN Cheng et al. (2021); (B) Graph Transformer Cai et al. (2022); (C) Recurrent CNN Lam et al. (2020); (D) SFCN Peng et al. (2021); (E) DGA-DMIL (GoogleNet backbone); (F) Our proposed DGA-DMIL framework (VGG backbone). In all figures, the green dashed line indicates the ideal estimation y = x (i.e., the estimated brain age equals the chronological age), and the red line is the linear regression model fitted by the estimated brain age and chronological age. MAE: mean absolute error; RMSE: root means squared error; PCC: Pearson correlation coefficient between the estimated brain age and chronological age.
@@ 5.1 Parameter Evaluation
Once we have obtained the bag of instances, we feed them into a backbone network to extract deep feature maps. We further evaluate the performance of our proposed method based on different CNN backbones, including GoogleNet, ResNet18, and VGG13. As shown in TABLE 3, the VGG13 backbone achieves the best performance in brain age estimation, which is consistent with the results of a similar study reported in He et al. (2021a). To verify the effectiveness of the proposed graph attention-based aggregator, we compare it with a CNN-based aggregator. The CNN aggregator uses two convolution layers and activity layers to reduce the channel of the feature maps. FC layers are then adopted to reduce the channel to 1 for score generation and feature aggregation. Additionally, we compare the effects of different backbones in combination with different MIL pooling strategies. As demonstrated in TABLE 3, the VGG backbone with graph attention aggregator outperforms all other competing combinations. Therefore, in our experiments, we use VGG as the CNN backbone and GAT as the aggregator.
@@ 5.2 Comparison with Other MIL Models of Brain Age Estimation
In this experiment, we test the performance of the proposed network on brain age estimation using the UKB dataset. We compare the proposed dual graph attention-based MIL method with other MIL methods, including Mean MIL, Deep MIL Ilse et al. (2018), DA-CMIL Chikontwe et al. , and TransMIL Shao et al. (2021). Mean MIL is a traditional MIL method that uses a mean operator for all MIL instances. In our experiments, we employ global mean pooling for the mean operator. Deep MIL adopts a weighted average of instances (low-dimensional embeddings) where the weights are determined by a neural network. DA-CMIL proposes attention-based MIL pooling and contrastive learning to better model the instance. TransMIL introduces a revised self-attention mechanism to reduce the computational complexity of MIL pooling. In addition, TransMIL employs convolution kernels of different sizes in the same layer to encode positional information. The above methods are implemented with our best efforts based on the published codes. These results are provided in TABLE 4. Our proposed method showed an improvement of 0.33 in the accuracy of brain age prediction. The performance of our proposed MIL model is consistently better than the compared methods.
@@
Mean MIL 2.86$\pm$0.19 3.55$\pm$0.24 0.9074$\pm$0.0068
Deep MILIlse et al. (2018) 2.64$\pm$0.18 3.30$\pm$0.21 0.9092$\pm$0.0066
DA-CMILChikontwe et al. 2.55$\pm$0.10 3.18$\pm$0.12 0.9278$\pm$0.0056
Trans-MILShao et al. (2021) 2.45$\pm$0.16 3.08$\pm$0.19 0.9138$\pm$0.0044
@@ 5.3 Comparison with SoTA Methods of Brain Age Estimation
We perform a comprehensive comparison of the proposed method with eight state-of-the-art approaches recently published in the literature, and evaluate them focusing on MAE, MSE, and PCC metrics. The methods compared mainly use deep neural networks, including CNNs, graph-based models, and transformers, for brain age estimation based on T1-weighted sMRI data. To ensure a fair comparison, all methods are implemented with the same training configuration and evaluated on identical training, validation, and test datasets. In this experiment, we implement the public source code of the methods  Cai et al. (2022); He et al. (2021a); Lam et al. (2020); Cheng et al. (2021); Peng et al. (2021) available on GitHub. However, for methods without a PyTorch implementation, such asBashyam et al. (2020); Lee et al. (2022); Yin et al. (2023), we made our best efforts to reproduce the methods based on the details provided in their respective papers. It’s worth noting that these 3D CNN models follow a standard architecture and are relatively easy to implement.
@@
Bashyam et al., 2020 Bashyam et al. (2020) 3.26$\pm$0.20 4.05$\pm$0.22 0.8604$\pm$0.0029
Lam et al., 2020 Lam et al. (2020) 2.41$\pm$0.11 3.06$\pm$0.14 0.9307$\pm$0.0012
He et al., 2021 He et al. (2021a) 3.10$\pm$0.23 3.85$\pm$0.28 0.8515$\pm$0.0161
Cheng et al., 2021 Cheng et al. (2021) 2.66$\pm$0.05 3.35$\pm$0.06 0.8982$\pm$0.0022
Peng et al., 2021 Peng et al. (2021) 2.34$\pm$0.10 2.93$\pm$0.12 0.9222$\pm$0.0016
Cai et al., 2022 Cai et al. (2022) 3.13$\pm$0.27 3.95$\pm$0.29 0.8582$\pm$0.0091
Lee et al., 2022 Lee et al. (2022) 2.61$\pm$0.12 3.25$\pm$0.14 0.9282$\pm$0.0019
Yin et al., 2023 Yin et al. (2023) 2.76$\pm$0.24 3.44$\pm$0.29 0.9076$\pm$0.0054
@@
He et al., 2021 He et al. (2021a)$\diamond$ 3.44$\pm$0.53 4.56$\pm$0.82 0.4169$\pm$0.2162
He et al., 2021 He et al. (2021a)$\ast$ 3.32$\pm$0.54 4.34$\pm$0.66 0.5527$\pm$0.0872
Peng et al., 2021 Peng et al. (2021$\diamond$ 3.51$\pm$0.55 4.52$\pm$0.67 0.4545$\pm$0.1489
Peng et al., 2021 Peng et al. (2021$\ast$ 3.52$\pm$0.57 4.48$\pm$0.57 0.4817$\pm$0.1228
Lee et al., 2022 Lee et al. (2022$\diamond$ 3.48$\pm$0.48 4.57$\pm$0.37 0.4799$\pm$0.1083
Lee et al., 2022 Lee et al. (2022$\ast$ 3.02$\pm$0.28 3.86$\pm$0.28 0.6536$\pm$0.0708
@@ 5.4 Ablation Experiments of the Framework
To evaluate the effect of the proposed dual graph attention aggregator and disentanglement branch in the proposed framework, we consider several settings where the dual graph attention aggregator and disentanglement branch are either employed or not (Figure 7 and TABLE 7). Specifically, if the dual graph attention aggregator is excluded, the framework would need to be modified in two aspects; (1) without instance-level MIL pooling of features, we default to using global average pooling (GAP) of instance features for simplicity (denoted as w/o aggregator $A_{\theta,S}$), and (2) without spatial feature aggregation in instances (spatial MIL pooling), one can opt to use the mean of instance features to obtain the overall bag-level feature z (denoted as w/o aggregator $A_{\theta,S}$, $A_{\theta,I}$). It is worth noting that the results without dual aggregator are consistent with the results with mean MIL pooling since we utilize GAP (equivalent to mean MIL pooling) instead of Graph Attention MIL pooling.
@@ 6 DISCUSSION
While our method has demonstrated good performance in estimating brain age on the UK Biobank and ADNI datasets, it is important to acknowledge the presence of certain limitations that should be addressed in future research. Firstly, our assumption that an individual’s actual age corresponds to the biological age of their brain may not always hold. To improve the accuracy of age prediction, future work could explore the integration of age-related information from multiple organs, such as the heart and blood vessels. Secondly, the primary objective of this study is to develop an accurate estimation model for brain MRI. It is worth noting that, similar to other studies in the literature He et al. (2021a, 2022); Peng et al. (2021), patient MRIs are not included in this research. Thirdly, our proposed method focuses primarily on analyzing brain sMRI data and does not incorporate text-based information such as gender, race, or ethnicity, which could potentially affect the manifestation of brain age effects. In future studies, it would be valuable to investigate the inclusion of text-based variables to better understand the differential effects of brain age across different demographic groups. This could involve building separate models or accounting for the influence of gender, race, ethnicity, and other relevant factors. Additionally, imputation techniques could be used to further refine our method. By using imputation techniques, we can potentially increase the completeness and accuracy of the data, leading to improved brain age estimation performance.

### v2

@@ 4.1 Datasets and Preprocessing
A standard preprocessing pipeline using FMRIB Software Library (FSL, https://fsl.fmrib. ox.ac.uk/[47, 45, 27] is applied to all sMRI data. It includes several steps such as nonlinear registration [2] to the MNI152 standard space, brain extraction [44, 28], and normalization of voxel values. To ensure consistency of input sizes across different datasets, all pre-processed sMRI images are cropped based on a generated mask that removes uninformative zero-value pixels from the background. As a result, all MRIs have a voxel size of $160\times 192\times 160$ and an isotropic spatial resolution of 1 $mm^{3}$.
@@ 4.3 Network Training
To train the proposed DGA-DMIL network, we use ADAM optimizer[31] for optimization, and initialize it with a learning rate of 0.0001. The batch size is set to be 32 and the disentanglement weights agree $\lambda_{2}=0.1$, $\lambda_{3}=0.05$, and $\lambda_{4}=1$, based on their performance on the validation set. The weights in the CNN layers in the decoupling unit are initialized using the Kaiming weight initialization scheme [19], and the weights in the fully connected layers are initialized with a normal distribution. The bias of the final fully connected layer is set equal to the average age of the training dataset for faster convergence. During training, the sum of the mean squared error loss and the decoupling loss is adopted as the total loss, and backpropagation is performed accordingly. All brain age estimation models are implemented in PyTorch.
@@
Fig. 6: Scatter diagrams of the estimated brain ages and chronological ages by different estimation models. (A) TSAN [8]; (B) Graph Transformer [5]; (C) Recurrent CNN [33]; (D) SFCN [41]; (E) DGA-DMIL (GoogleNet backbone); (F) Our proposed DGA-DMIL framework (VGG backbone). In all figures, the green dashed line indicates the ideal estimation y = x (i.e., the estimated brain age equals the chronological age), and the red line is the linear regression model fitted by the estimated brain age and chronological age. MAE: mean absolute error; RMSE: root means squared error; PCC: Pearson correlation coefficient between the estimated brain age and chronological age.
@@ 5.1 Parameter Evaluation
Once we have obtained the bag of instances, we feed them into a backbone network to extract deep feature maps. We further evaluate the performance of our proposed method based on different CNN backbones, including GoogleNet, ResNet18, and VGG13. As shown in TABLE 3, the VGG13 backbone achieves the best performance in brain age estimation, which is consistent with the results of a similar study reported in [21]. To verify the effectiveness of the proposed graph attention-based aggregator, we compare it with a CNN-based aggregator. The CNN aggregator uses two convolution layers and activity layers to reduce the channel of the feature maps. FC layers are then adopted to reduce the channel to 1 for score generation and feature aggregation. Additionally, we compare the effects of different backbones in combination with different MIL pooling strategies. As demonstrated in TABLE 3, the VGG backbone with graph attention aggregator outperforms all other competing combinations. Therefore, in our experiments, we use VGG as the CNN backbone and GAT as the aggregator.
@@ 5.2 Comparison with Other MIL Models of Brain Age Estimation
In this experiment, we test the performance of the proposed network on brain age estimation using the UKB dataset. We compare the proposed dual graph attention-based MIL method with other MIL methods, including Deep MIL [26], DA-CMIL [10], and TransMIL [42]. In our experiments, we employ global mean pooling for the mean operator. Deep MIL adopts a weighted average of instances (low-dimensional embeddings) where the weights are determined by a neural network. DA-CMIL proposes attention-based MIL pooling and contrastive learning to better model the instance. TransMIL introduces a revised self-attention mechanism to reduce the computational complexity of MIL pooling. In addition, TransMIL employs convolution kernels of different sizes in the same layer to encode positional information. The above methods are implemented with our best efforts based on the published codes. These results are provided in TABLE 4. Our proposed method showed an improvement of 0.33 in the accuracy of brain age prediction. The performance of our proposed MIL model is consistently better than the compared methods.
@@
Deep MIL[26] 2.64$\pm$0.18 3.30$\pm$0.21 0.9092$\pm$0.0066
DA-CMIL[10] 2.55$\pm$0.10 3.18$\pm$0.12 0.9278$\pm$0.0056
Trans-MIL[42] 2.45$\pm$0.16 3.08$\pm$0.19 0.9138$\pm$0.0044
@@ 5.3 Comparison with SoTA Methods of Brain Age Estimation
We perform a comprehensive comparison of the proposed method with eight state-of-the-art approaches recently published in the literature, and evaluate them focusing on MAE, MSE, and PCC metrics. The methods compared mainly use deep neural networks, including CNNs, graph-based models, and transformers, for brain age estimation based on T1-weighted sMRI data. To ensure a fair comparison, all methods are implemented with the same training configuration and evaluated on identical training, validation, and test datasets. In this experiment, we implement the public source code of the methods  [5, 21, 33, 8, 41] available on GitHub. However, for methods without a PyTorch implementation, such as[3, 34, 49], we made our best efforts to reproduce the methods based on the details provided in their respective papers. It’s worth noting that these 3D CNN models follow a standard architecture and are relatively easy to implement.
@@
Bashyam et al., 2020 [3] 3.26$\pm$0.20 4.05$\pm$0.22 0.8604$\pm$0.0029
Lam et al., 2020 [33] 2.41$\pm$0.11 3.06$\pm$0.14 0.9307$\pm$0.0012
He et al., 2021 [21] 3.10$\pm$0.23 3.85$\pm$0.28 0.8515$\pm$0.0161
Cheng et al., 2021 [8] 2.66$\pm$0.05 3.35$\pm$0.06 0.8982$\pm$0.0022
Peng et al., 2021 [41] 2.34$\pm$0.10 2.93$\pm$0.12 0.9222$\pm$0.0016
Cai et al., 2022 [5] 3.13$\pm$0.27 3.95$\pm$0.29 0.8582$\pm$0.0091
Lee et al., 2022 [34] 2.61$\pm$0.12 3.25$\pm$0.14 0.9282$\pm$0.0019
Yin et al., 2023 [49] 2.76$\pm$0.24 3.44$\pm$0.29 0.9076$\pm$0.0054
@@
He et al., 2021 [21$\diamond$ 3.44$\pm$0.53 4.56$\pm$0.82 0.4169$\pm$0.2162
He et al., 2021 [21$\ast$ 3.32$\pm$0.54 4.34$\pm$0.66 0.5527$\pm$0.0872
Peng et al., 2021 [41$\diamond$ 3.51$\pm$0.55 4.52$\pm$0.67 0.4545$\pm$0.1489
Peng et al., 2021 [41$\ast$ 3.52$\pm$0.57 4.48$\pm$0.57 0.4817$\pm$0.1228
Lee et al., 2022 [34$\diamond$ 3.48$\pm$0.48 4.57$\pm$0.37 0.4799$\pm$0.1083
Lee et al., 2022 [34$\ast$ 3.02$\pm$0.28 3.86$\pm$0.28 0.6536$\pm$0.0708
@@ 5.4 Ablation Experiments of the Framework
To evaluate the effect of the proposed dual graph attention aggregator and disentanglement branch in the proposed framework, we consider several settings where the dual graph attention aggregator and disentanglement branch are either employed or not (Figure 7 and TABLE 7). Specifically, if the dual graph attention aggregator is excluded, the framework would need to be modified in two aspects; (1) without instance-level MIL pooling of features, we default to using global average pooling (GAP) of instance features for simplicity (denoted as w/o aggregator $A_{\theta,S}$), and (2) without spatial feature aggregation in instances (spatial MIL pooling), one can opt to use the mean of instance features to obtain the overall bag-level feature z (denoted as w/o aggregator $A_{\theta,S}$, $A_{\theta,I}$). It is worth noting that the results without dual aggregator are consistent with the results with mean MIL pooling since we utilize GAP instead of Graph Attention MIL pooling.
@@ 6 DISCUSSION
While our method has demonstrated good performance in estimating brain age on the UK Biobank and ADNI datasets, it is important to acknowledge the presence of certain limitations that should be addressed in future research. Firstly, our assumption that an individual’s actual age corresponds to the biological age of their brain may not always hold. To improve the accuracy of age prediction, future work could explore the integration of age-related information from multiple organs, such as the heart and blood vessels. Secondly, the primary objective of this study is to develop an accurate estimation model for brain MRI. It is worth noting that, similar to other studies in the literature [21, 20, 41], patient MRIs are not included in this research. Thirdly, our proposed method focuses primarily on analyzing brain sMRI data and does not incorporate text-based information such as gender, race, or ethnicity, which could potentially affect the manifestation of brain age effects. In future studies, it would be valuable to investigate the inclusion of text-based variables to better understand the differential effects of brain age across different demographic groups. This could involve building separate models or accounting for the influence of gender, race, ethnicity, and other relevant factors. Additionally, imputation techniques could be used to further refine our method. By using imputation techniques, we can potentially increase the completeness and accuracy of the data, leading to improved brain age estimation performance.
@@ 7 CONCLUSION
ACKNOWLEDGMENTS
This research has been conducted using the UK Biobank Resource under Application Number 94178.